# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v1

@@ 4.2. General clone is not stronger than blanket
For every local randomizer, there is a post-processing function from $(P_{0}^{GC},P_{1}^{GC})$ to the $(P_{0}^{B},P_{1}^{B})$, where $(P_{0}^{GC},P_{1}^{GC})$ is the reduction pair given by the general clone paradigm equipped with any decomposition, and $(P_{0}^{B},P_{1}^{B})$ is the reduction pair given by the privacy blanket framework. Therefore,
$D_{e^{\epsilon}}(\mathcal{M}_{S}(X_{0}),\mathcal{M}_{S}(X_{1}))\leq D_{e^{\epsilon}}(P_{0}^{B},P_{1}^{B})\leq D_{e^{\epsilon}}(P_{0}^{GC},P_{1}^{GC}).$
@@
The core idea of the proof is that the privacy blanket characterizes the maximal common part shared by all $\mathcal{R}(x_{i})$. The common part in any decomposition under the general clone paradigm cannot exceed that of the privacy blanket.
Recall the definition of $\omega(y)=\inf_{x\in\mathbb{X}}R(x)(y)/\gamma$, where $\gamma$ is a normalization factor. An important observation is that
$\forall y\in\mathbb{Y}:\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)\leq\gamma\omega(y).$
It is because the formula (2) should always hold for all $x_{2},x_{3},\dots,x_{n}\in\mathbb{X}$, i.e., $\forall x\in\mathbb{X}:\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)\leq R(x)(y)$. Hence, it follows that $\sum_{j=1}^{k}c_{j}\leq\gamma$ and $1-\gamma\leq\beta$. It means that each $Q_{i}$ can be decomposed by
$Q_{i}=\frac{\gamma-\sum_{j=1}^{k}c_{j}}{\beta}Q^{com}+(1-\frac{\gamma-\sum_{j=1}^{k}c_{j}}{\beta})Q_{i}^{\prime},$
where $Q^{com},Q_{i}^{\prime}$ are two probability distributions, $Q^{com}$ is the common part of all $Q_{i}$.
The function $f_{2}$ shown in Algorithm 3 is the post-processing function satisfying $f_{2}(P_{0}^{GC})=P_{0}^{B},f_{2}(P_{1}^{GC})=P_{1}^{B}$. It behaves as follows: when encountering an index $i\in[1,k]$, it samples from the corresponding distribution $Q_{1}^{i}$; when encountering $i=k+1$, it samples from the common distribution $Q^{\text{com}}$ with a certain probability. Taken together, this behavior is equivalent to each user (except the first) sampling from the blanket distribution with probability $\gamma$. ∎
As a result, although both $D_{e^{\epsilon}}(P_{0}^{GC}||P_{1}^{GC})$ and $D_{e^{\epsilon}}(P_{0}^{B}||P_{1}^{B})$ serve as upper bounds on the privacy amplification in the shuffle model, the bound provided by the privacy blanket is always at least as tight as that of the general clone paradigm under any decomposition.
4.3. Blanket is “in” the General Clone
Theorem 4.
For every local randomizer, the following decomposition in the general clone paradigm is equivalent to the privacy blanket framework:
$\mathcal{R}(x_{1}^{0})$ $=\sum_{j=1}^{|\mathbb{Y}|}a_{j}\mathbbm{1}_{y_{j}}$
$\mathcal{R}(x_{1}^{1})$ $=\sum_{j=1}^{|\mathbb{Y}|}b_{j}\mathbbm{1}_{y_{j}}$
(3) $\forall i\in[2,n],\quad\mathcal{R}(x_{i})$ $=\sum_{j=1}^{|\mathbb{Y}|}c_{j}\mathbbm{1}_{y_{j}}+\beta Q_{i}.$
Proof.
It is straightforward to observe that $(P_{0}^{B},P_{1}^{B})$ and $(P_{0}^{GC},P_{1}^{GC})$ are essentially equivalent, differing only in some technical notations. ∎
We refer to the optimal decomposition of a local randomizer as the decomposition equivalent to that in the privacy blanket framework. For simplicity, we can merge some components to obtain an equally optimal decomposition. As an example, consider the $k$-Random Response mechanism.
Example 4.0.
Its optimal decomposition is as follows:
$R(x_{1}^{0})$ $=e^{\epsilon_{0}}p\mathbbm{1}_{x_{1}^{0}}+p\mathbbm{1}_{x_{1}^{1}}+qU,$
$R(x_{1}^{1})$ $=p\mathbbm{1}_{x_{1}^{0}}+e^{\epsilon_{0}}p\mathbbm{1}_{x_{1}^{1}}+qU,$
$\forall i\in[2,n]:R(x_{i})$ $=p\mathbbm{1}_{x_{1}^{0}}+p\mathbbm{1}_{x_{1}^{1}}+qU+(e^{\epsilon_{0}}-1)p\mathbbm{1}_{x_{i}}.$
where $p=\frac{1}{e^{\epsilon_{0}}+k-1}$, $q=(k-2)p$, and $U$ is the uniform distribution over $[k]-\{x_{1}^{0},x_{1}^{1}\}$. This decomposition is also considered in the corrected version of the stronger clone (Feldman et al., 2023b).
@@ 4.3. Blanket is “in” the General Clone
Algorithm 3 Post-processing function from general clone paradigm to privacy blanket framework, $f_{2}$
$(n_{1},\dots,n_{k},n_{k+1})\in\{0,1,\dots,n\}^{k}$
@@
for $i=1,\dots,k+1$ do
     for $count=1,2,\dots,n_{i}$ do
         if $i=k+1$ then
              $r\leftarrow\text{Bern}(\frac{\gamma-\sum_{j=1}^{k}c_{j}}{\beta})$
@@
                  $s\leftarrow_{Q^{com}}\mathbb{Y}$triangleright Sample from $Q^{com}$
@@
              $s\leftarrow_{Q^{i}_{1}}\mathbb{Y}$triangleright Sample from $Q^{i}_{1}$

### v5

@@ 4.3 General clone is not stronger than blanket
For every local randomizer, there is a post-processing function from $(P_{0}^{\mathsf{GC}},P_{1}^{\mathsf{GC}})$ to the $(P_{0}^{\mathsf{B}},P_{1}^{\mathsf{B}})$, where $(P_{0}^{\mathsf{GC}},P_{1}^{\mathsf{GC}})$ is the reduction pair given by the general clone paradigm equipped with any decomposition, and $(P_{0}^{\mathsf{B}},P_{1}^{\mathsf{B}})$ is the reduction pair given by the privacy blanket framework. Therefore,
$D_{e^{\varepsilon}}(\mathcal{M}_{S}(X_{0})\parallel\mathcal{M}_{S}(X_{1}))\leq D_{e^{\varepsilon}}(P_{0}^{\mathsf{B}}\parallel P_{1}^{\mathsf{B}})\leq D_{e^{\varepsilon}}(P_{0}^{\mathsf{GC}}\parallel P_{1}^{\mathsf{GC}}).$
@@
The core idea is that the privacy blanket characterizes the maximal common component shared by all $\mathcal{R}(x_{i})$, and no decomposition under the general clone paradigm can yield a larger common part. Recall that the privacy blanket is defined via:
$\omega(y)=\frac{1}{\gamma}\inf_{x\in\mathbb{X}}\mathcal{R}(x)(y),$
where $\gamma$ is a normalization constant. An important observation is:
$\forall y\in\mathbb{Y}:\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)\leq\gamma\omega(y).$ (4)
This holds because the decomposition condition (2) in Definition 7 must be satisfied for all inputs $x_{2},x_{3},\dots,x_{n}\in\mathbb{X}$, i.e.,
$\forall x\in\mathbb{X}:\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)\leq\mathcal{R}(x)(y).$
Taking the infimum over $x\in\mathbb{X}$ on both sides of the above inequality yields (4). Integrating both sides of (4) over $y$ yields $\sum_{j=1}^{k}c_{j}\leq\gamma$.
Define a distribution $Q^{\text{com}}$ as the common part of all $Q_{i}$:
$Q^{\text{com}}(y)=\frac{\gamma\omega(y)-\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)}{\mu},$
where $\mu=\gamma-\sum_{j=1}^{k}c_{j}$ is a normalization factor. When $\mu=0$, the distribution $Q^{\mathrm{com}}$ can be defined arbitrarily. Then, each $Q_{i}$ can be decomposed by
$Q_{i}=\frac{\mu}{\beta}Q^{\text{com}}+(1-\frac{\mu}{\beta})Q_{i}^{\prime}.$
Here, $Q_{i}^{\prime}$ is a probability distribution, whose existence is established as follows: From equation (2) and the definition of $\omega$, we have:
$\forall i\in[2,n],\forall y\in\mathbb{Y}:\gamma\omega(y)\leq\mathcal{R}(x_{i})(y)=\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)+\beta Q_{i}(y).$
This implies:
$\forall i\in[2,n],\forall y\in\mathbb{Y}:Q_{i}(y)\geq\frac{\gamma\omega(y)-\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)}{\beta}=\frac{\mu}{\beta}Q^{\text{com}}(y).$
Let $f_{3}$ be the post-processing function defined in Algorithm 5. Its behavior is as follows: it processes each input sequentially. When the input is $j\in[1,k]$, it outputs a sample from $Q_{1}^{j}$; when $j=k+1$, it samples from $Q^{\text{com}}$ with probability $\mu/\beta$.
We now argue that $f_{3}(P_{0}^{\mathsf{GC}})=P_{0}^{\mathsf{B}}$. By definition, $P_{0}^{\mathsf{B}}$ outputs one sample from $\mathcal{R}(x_{1}^{0})$, and $|\boldsymbol{b}|\sim\text{Bin}(n-1,\gamma)$ samples from $\omega$. From Definition 8, $P_{0}^{\mathsf{GC}}$ draws one sample from $A_{0}$, and $n-1$ samples from $A_{2}$. Since $\text{supp}(A_{0})\subseteq\{1,\dots,k\}$, its output is always retained. The transformed output has distribution:
$\Pr[f_{3}(A_{0})=y]$ $=\sum_{j=1}^{k}\Pr[A_{0}=j]\cdot\Pr[f_{3}(j)=y]$
$=\sum_{j=1}^{k}a_{j}Q_{1}^{j}(y)=\mathcal{R}(x_{1}^{0})(y).$
Now consider $A_{2}$. The event $f_{3}(A_{2})=\emptyset$ occurs with probability:
$\Pr[f_{3}(A_{2})=\emptyset]$ $=\Pr[A_{2}=k+1]\cdot\Pr[f_{3}(k+1)=\emptyset]$
$=\beta\cdot\left(1-\frac{\mu}{\beta}\right)=\beta+\sum_{j=1}^{k}c_{j}-\gamma=1-\gamma.$
Moreover, the conditional distribution of $f_{3}(A_{2})$, given that it is non-empty, is exactly $\omega$:
$\Pr[f_{3}(A_{2})=y|f_{3}(A_{2})\neq\emptyset]$
$=\sum_{j=1}^{k+1}\Pr[A_{2}=j|f_{3}(A_{2})\neq\emptyset]\cdot\Pr[f_{3}(j)=y|f_{3}(A_{2})\neq\emptyset\wedge A_{2}=j]$
$=\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)/\gamma+\frac{\mu}{\gamma}\cdot Q^{\text{com}}(y)$
$=\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)/\gamma+\frac{\mu}{\gamma}\cdot\frac{\gamma\omega(y)-\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)}{\mu}$
$=\omega(y).$
Hence, $f_{3}(P_{0}^{\mathsf{GC}})=P_{0}^{\mathsf{B}}$, and the same argument applies for $f_{3}(P_{1}^{\mathsf{GC}})=P_{1}^{\mathsf{B}}$. This completes the proof. ∎
@@
Algorithm 5 Post-processing function from general clone to privacy blanket framework, $f_{3}$
$(n_{1},\dots,n_{k},n_{k+1})\in\{0,1,\dots,n\}^{k+1}$
@@
for $j=1,\dots,k+1$ do
     for $count=1,2,\dots,n_{j}$ do
         if $j=k+1$ then
              $r\leftarrow\text{Bern}(\frac{\gamma-\sum_{m=1}^{k}c_{m}}{\beta})$
@@
                  $s\leftarrow_{Q^{\text{com}}}\mathbb{Y}$triangleright Sample from $Q^{\text{com}}$
@@
              $s\leftarrow_{Q^{j}_{1}}\mathbb{Y}$triangleright Sample from $Q^{j}_{1}$